OpenStudy (anonymous):

How is Y=2^x geometrically transformed into 3^5x/2

OpenStudy (anonymous):

Is that \[y=2^x, y=3^{\frac{5x}{2}}\]?

OpenStudy (anonymous):

Yes, it is

OpenStudy (anonymous):

I thought it was a stretch in the x-axis by 2.5, Apparently, it's by 0.4??

OpenStudy (anonymous):

Ohhh I made a mistake..! It's supposed to be\[ Y=3^x\]

OpenStudy (anonymous):

and another one like that is the transformation of \[y=3^x\] onto \[y=9^x\] Again, I got a stretch in the x-axis by S.F. 3, but it's by S.F 0.5

OpenStudy (anonymous):

let look at \[y=ab^{cx}\] okay it seems that to find the stretch in the x axis it will be \[\frac{1}{c}\] I think that solves the problem. I'll keep looking to find a site that can show this with a general form.

OpenStudy (anonymous):

For the second example you wrote the equations will be \[y=3^x, y=9^x=3^{2x}\] So the stretch is 1/2 or 0.5

OpenStudy (anonymous):

I see what you mean, but I'm not sure why that's the formula lol...

OpenStudy (anonymous):

This site explains it well the only thing is that it looks at e^x graphs which aren't any different from your ones. http://archives.math.utk.edu/visual.calculus/0/shifting.5/index.html

OpenStudy (anonymous):

Oh, Ok. Thanks :)

OpenStudy (anonymous):

If you type both equations into Wolfram Alpha you can see it visually as well. For y=3^x from -2 to 2 is shown exactly the same as y=9^x but only for -1 to 1. So half the space. I'm still not sure on the formula though, sorry.